# Learning in Chaos: Efficient Autoscaling and Self-healing for Distributed Training at the Edge

### v1

@@ Abstract
Frequent node and link changes in edge AI clusters disrupt distributed training, while traditional checkpoint-based recovery and cloud-centric autoscaling are too slow for scale-out and ill-suited to chaotic and self-governed edge. This paper proposes Chaos, a resilient and scalable edge distributed training system with built-in self-healing and autoscaling. It speeds up scale-out by using multi-neighbor replication with fast shard scheduling, allowing a new node to pull the latest training state from nearby neighbors in parallel while balancing the traffic load between them. It also uses a cluster monitor to track resource and topology changes to assist scheduler decisions, and handles scaling events through peer negotiation protocols, enabling fully self-governed autoscaling without a central admin. Extensive experiments show that Chaos consistently achieves much lower scale-out delays than Pollux, EDL, and Autoscaling, and handles scale-in, connect-link, and disconnect-link events within 1 millisecond, making it smoother to handle node joins, exits, and failures. It also delivers the lowest idle time, showing superior resource use and scalability as the cluster grows.
@@
Distributed machine learning, autoscaling, self-healing, elastic training.
@@ I Introduction
The rapid growth of model and cluster sizes is pushing AI training to its limits. At OpenAI, a hidden bug caused frequent crashes during GPT-4.5 training, forcing constant checkpointing and restarts [1]. DeepSeek [2] reallocates inference machines for model training at night, requiring clusters to scale up and down on the fly. The demand for self-healing and autoscaling is even more urgent in edge AI training, where devices are more unstable and free to join and leave.
@@
    Edge devices and their interconnections are more prone to interruptions. They are decentralized and unpredictable, they can drop out at any time due to failures in network, device and software, and power loss, movement, user interrupts. Even if devices remain alive, unstable connections caused by routing failures, firewalls, congestion, or mobility can still disrupt training.
@@
    Edge devices are self-governed and free to join, leave, and switch networks at will. They may join when idle or charging, and leave under heavy load or to save RAM/CPU/energy. They are also free to join or leave a network, leading to new or terminated connections.
@@
Changes in cluster nodes and links can interrupt distributed training. A common fix to handle unexpected node dropouts is checkpointing and restart: save model checkpoints, fix the issue, then roll back and restart training from checkpoints [1, 3, 4, 5, 6, 7]. This stop-resume method also applies to autoscaling: stop training when nodes join or leave, reinit the cluster, and restart training from the last checkpoint [8, 9, 10, 11, 12]. This process is time- and resource-costly due to heavy disk I/O from frequent checkpointing. To address this, stop-free autoscaling [13, 14, 15, 16, 17] lets new nodes pull the latest training state from running nodes, no restarts, and no checkpoints. This enables continuous and scalable training at chaotic edge.
While these systems have supported self-healing or autoscaling, they are slow in practice. When a new node joins, it must pull the latest training state, such as model weights and optimizer state, from a running node. This time cost is called state replication delay. In data centers, fast and symmetric LANs keep this delay low, even when pulling from a random node [8, 9, 10, 11, 13, 14, 15, 16, 17]. At the edge, however, limited bandwidth can stretch delays to tens of seconds, or even minutes if the pulling node is far away in an asymmetric, uneven network. Such high delays are unacceptable for fast-changing edge clusters. To keep training smooth, replication must be optimized for fast act to node and link changes.
Secondly, existing systems are designed for the HPC, not the edge. They support either self-healing or autoscaling, but edge clusters demand both because the network environment is far more chaotic. In addition, current autoscaling is centrally managed, i.e., there is an admin monitoring cluster load and manually triggering scaling up/down. In contrast, edge devices are decentralized and self-governed, with no central control. As a result, edge training systems must react on their own, detecting node and link changes and fixing communication rules at runtime with no admin involved.
This paper proposes Chaos, a resilient and scalable edge distributed training system with self-healing and autoscaling. Chaos monitors cluster and resource changes in real time. When nodes or links join, leave, or fail, it automatically triggers scale-in, scale-out, connect-link, and disconnect-link actions. It removes faulty nodes or links, resolves cluster blocking, scales the cluster up or down, and fixes communication rules on the fly. To speed up state replication for new nodes, we introduce a neighbor-aware mechanism. It uses the overlay network topology and resources detected by Chaos to find neighbor nodes, and pulls the training state from them in parallel, using multi-path transfer and traffic scheduling for faster replication. We set up an edge cluster with 6–12 virtual machines, simulating node joins and exits under constrained, heterogeneous network conditions. The results show that, compared to alternatives in Pollux [8], EDL [13], and Autoscaling [18], our Chaos handles self-healing and autoscaling more efficiently.
Our main contributions are summarized as follows:
@@
  • We propose a resilient, scalable edge distributed training system with self-healing and autoscaling that can more efficiently handles node and link joins, exits, and failures.
  • We propose a multi-neighbor state replication mechanism that pulls the training state from neighbors in parallel, with a shard scheduler to balance their traffic load.
@@
  • We propose a cluster monitor that tracks resource and topology changes, along with peer negotiation protocols that enable autoscaling without a central admin.
@@
  • We implement Chaos and show that it achieves much lower scale-out delays than Pollux, EDL, and Autoscaling, with superior resource efficiency and scalability.
@@ II-A Reviews on Self-healing and Autoscaling
Self-healing. In distributed machine learning (DML), worker nodes frequently synchronize model weights, but in edge networks, frequent node and link failures block synchronization and freeze training. Checkpointing helps recover from such failures by periodically saving the training state, e.g., model weights, optimizer states, so the system can restart from the last checkpoint [1, 7]. DataStates-LLM [3] and CheckFreq [19] reduce training stalls by asynchronously writing checkpoints. JIT [4] triggers checkpointing on other nodes only after a failure occurs. Check-N-Run [5] uses differential storage to cut checkpoint size. Gemini [20] speed up recovery by redundantly storing checkpoints in RAM across devices.
@@
Autoscaling. Checkpointing is also a key approach to autoscaling, known as the “stop-resume” strategy [13]. It saves a checkpoint, pauses training, resizes the cluster, then restarts training from the latest checkpoint. To reduce the checkpointing overhead, Optimus [9] triggers autoscaling detection every ten minutes. Pollux [8] saves checkpoints only when cluster size and topology need to change. Rubick [12] sets an empirical threshold and only reconfigures when restart overhead stays below it.
@@
Stop-free autoscaling offers a faster and more flexible alternative. When a new node joins, it pulls the training state directly from active nodes’ memory, so no need to restart training or load/write checkpoints from/to a slow disk. In parameter server setups [21], PaddleEDL [22] and ElasticDL [23] use asynchronous training to realize stop-free autoscaling. As workers run independently, their join and exit don’t disrupt others, and new nodes can simply fetch the latest model from the parameter server. However, they fall back to stop-resume in synchronous mode. DL2 [11] enables stop-free autoscaling in synchronous training using a scaling clock that detects new nodes and sends them the current model. In more general cases, systems like EDL [13], EPS [15], CoDDL [16], and DeepBoot [17] typically select a node (e.g., master) to pull the training state from. Elan [14] instead picks the neighbor with the highest bandwidth. Autoscaling [18] takes it further by slicing the training state across multiple nodes and pulling from them in parallel for faster replication.
A key limitation of these systems is that they are designed for centralized HPC clusters, not for the decentralized and autonomous edge. For example, Autoscaling adds nodes every few minibatches. EDL, EPS, Elan, CoDDL, and DeepBoot scale based on cluster load. DeepSeek offloads inference nodes to training clusters during idle hours. Pollux and Rubick schedule GPU resources and decide when and how much to scale. In these systems, a central admin makes the scaling decisions, and scaling happens in batches. However, the join and leave models at the edge are fundamentally different:
  • Chaos doesn’t need GPU scheduling like others do. User devices are free to join or leave at any time. Even if the admin could make optimal scheduling decisions, it cannot command users to join, as users may not cooperate.
  • Users’ joins and leaves follow a Poisson process, arriving one by one. This process follows the laws of nature, making GPU scheduling useless. Chaos must handle user joins and leaves one by one, unlike HPC schedulers that plan for batched changes.
Since user joins and leaves must be handled one by one, what matters more is how to make scaling fast, as scale-in and scale-out happen far more frequently than a batched change. This calls for a new, edge-native solution.
II-B Motivations
Question A: Which node(s) should a new node pull from for faster state replication?
Most systems let a new node pull the training state from a single node [13, 14, 15, 16, 17]. In the case shown in Figs. 1a and 1b, a new node 11 joins and picks the fastest neighbor to pull from, but sending the entire state over a single link creates a bottleneck, replication still takes 80 time units. Autoscaling [18] provides a multi-source solution by splitting the training state and pulling shards from multiple nodes in parallel. While effective in symmetric networks like fat-trees, it performs poorly in irregular edge networks, e.g., in Fig. 1c, where distant nodes require multi-hop forwards, causing delays of up to 88 time units, even along the shortest routes.
@@ II-B Motivations
Figure 1: An example of (a) the overlay network topology, (b) single-node replication, (c) multi-node replication, and (d) multi-neighbor replication. (e) compares their state replication delays. The number on each edge indicates the delay incurred to transfer one unit of data through that link. Suppose there are 10 data units to transmit.
It is clear that multi-hop forwards involves significant redundant transfers. In synchronous training, all nodes share the same model weights, so paths like 7→10→3→11 are wasteful, node 3 already has the data shards of nodes 7 and 10. Instead, node 3 alone can serve node 11, cutting unnecessary hops and transmission. This inspires multi-neighbor replication, where the new node pulls shards directly from its neighbors, as shown in Fig. 1d. With smart shard sizing, replication delay drops to 62 time units. This calls for a shard scheduler that minimizes the completion time to pull from all neighbors.
Question B: How does a new node join a decentralized, self-governed edge cluster?
Existing systems [18, 13, 14, 15, 16, 17] rely on an admin to manually trigger autoscaling. As shown in Fig. 2a, the new node keeps waiting until the admin decides to scale out and sends a command to initialize and join the cluster. This, however, does not work in a decentralized, self-governed edge cluster. There is no central admin to send such commands. Without it, new nodes are left waiting indefinitely. This motivates a shift: instead of waiting for a command that will never come, new nodes should take the initiative and actively send join requests. As shown in Fig. 2b, the new node actively connects to the cluster scheduler and runs stop-free autoscaling through self-discovery, peer negotiation, and state synchronization. Therefore, we need a cluster monitor to detect node and link changes, and peer-negotiation protocols for flexible scale-in, scale-out, connect-link, and disconnect-link events.
Figure 2: Illustration of node joins with and without an admin.
@@
To address question A, Section III mathematically models the shard assignment problem, introduces a polynomial-time algorithm, and analyzes its optimality gap and complexity. Then, for question B, Section IV presents the cluster monitor and peer negotiation protocols for scale-out, scale-in, connect-link, and disconnect-link. Section V outlines the overall system design, and Section VI provides the experiment results.

### v2

# Learning In Chaos: Efficient Autoscaling and Self-Healing for Multi-Party Distributed Training

@@ Abstract
Node and link churn in multi-party, cross-region clusters over wide-area networks (WANs) often disrupts distributed training. However, checkpoint-based recovery and cloud-centric autoscaling react slowly and assume centralized control, which is misaligned with the self-governed setup where institutions can freely join and leave. This paper proposes Chaos, a multi-party distributed training system with self-healing and autoscaling, enabling robust and elastic training under churn. It speeds up autoscaling via multi-neighbor state replication and model sharding. We formalize the sharding and assignment as a MINLP that captures WAN heterogeneity, and reduce it to a tractable MILP by analyzing its monotonicity on a divisibility chain. By establishing an equivalence, we derive a greedy algorithm that follows optimality rules and yields the optimal solution in polynomial time. Chaos uses a cluster monitor to track resource and topology changes, and handles scaling events through peer negotiation protocols, enabling fully self-governed autoscaling among institutions. Experiments show that Chaos has substantially lower scale-out delay than Pollux, Elan, and Autoscaling, and handles scale-in, connect-link, and disconnect-link events within 20ms. It also delivers the lowest idle time, showing superior resource use and scalability as the cluster grows.
@@
Distributed machine learning, dependable and fault-tolerant systems and networks, cross-region networks, network churn, elastic training, traffic balancing, peer protocols.
@@ I Introduction
Vast and critical datasets lie fragmented across individual institutions, creating barriers to innovation, especially in the medical and biotechnology industries. Cross-institution collaboration has emerged as an effective path to overcome these silos. In 2020, NVIDIA launched a federated learning (FL) project across 20 hospitals worldwide and developed an AI system to predict whether a patient with COVID-19 would require supplemental oxygen [1]. ADOPS used FL across breast imaging reports from 7 clinical institutions worldwide and developed a system for breast density classification [2]. In 2021, NVIDIA, UCLA, SUNY Upstate, and the NCI collaborated to develop an accurate prostate cancer segmentation model [3]. In 2022, NVIDIA and several U.S. universities used data from 42 hospitals across 5 healthcare systems in the U.S. and Europe to improve chest X-ray–based AI models for COVID-19 diagnosis [4]. Intel and Penn Medicine coordinated a federation of 71 medical and research institutions worldwide, achieving significant gains in brain tumor recognition [5]. In 2023, Owkin, NVIDIA, KCL, and over a dozen pharmaceutical partners used FL to develop an AI model on the world’s largest collaborative drug compound dataset [6]. Even in the new era of large models, collaborating multi-party data to apply better parameter-efficient fine-tuning on foundation models remains a major focus [7, 8, 9, 10, 11, 12].
These successful cases [1, 2, 3, 4, 5, 6] rely on heavy offline negotiations, where agreements must be signed in person and all institutions must synchronize to start a training task. This rigid process limits generalization: once joined, institutions cannot leave, and outsiders cannot join. To move beyond these constraints, a public and online coordination platform is essential. Such a platform should allow institutions to join or leave freely, as illustrated in Fig. 1. However, this openness and flexibility also raise new technical demands: the distributed system must support efficient self-healing and autoscaling to sustain robust training.
@@
    Need for self-healing: Institutions may go offline due to occasional device, power, network, or software failures. If the system cannot automatically detect the issue, reconfigure, and resume training, the entire system will fail.
@@
    Need for autoscaling: Institutions can participate in training during off-peak business hours and withdraw during peak hours or system maintenance. Without scaling out, new institutions cannot join; without scaling in, existing participants cannot leave.
Figure 1: A multi-party cluster under churn: (a) node joining, (b) node leaving, (c) link establishment, and (d) link loss.
@@
We abstract each institution as a node connected by WAN links, forming a logical overlay. Node or link churn alters this overlay’s topology and capacity, and interrupts distributed training. A common fix to handle unexpected node dropouts is checkpoint-restart: save model checkpoints, fix the issue, then roll back and restart training from checkpoints [13, 14, 15, 16, 17, 18]. This stop-resume method also applies to autoscaling: stop training when nodes join or leave, reinit the cluster, and restart training from the last checkpoint [19, 20, 21, 22, 23]. This process is time- and resource-costly due to heavy disk I/O or network transfer from frequent checkpointing. For example, checkpoint-restart takes $\sim$2min in Pollux [19] and $\sim$1min in Rubick [20] and UCP [23], preventing them from frequent scaling in/out. To address this, stop-free autoscaling [24, 25, 26, 27, 28, 29, 30, 31] lets new nodes synchronize training states from running nodes, without restarts or checkpoints. This enables continuous training over cross-region clusters with churn. However, they are still slow in practice. When a new node joins, it must fetch the latest training state from existing nodes, such as model weights. This time cost is referred to as state replication delay. In data centers, high-speed LANs keep this delay minimal. Across WANs, however, limited bandwidth can stretch the delay to several minutes, especially when the new node is geographically distant. Such high delays make autoscaling impractical for high-churn clusters. Thus, state replication should be optimized to adapt to node and link changes rapidly.
It should be noted that, unlike GPU scheduling systems in HPC (e.g., Pollux [19], Rubick [20], and [21, 22, 24, 25, 26, 27, 28, 29, 30, 31, 23]), which have global control over the cluster and can decide to add or remove nodes/GPUs during training for proactive scaling, our setting targets multiple self-governed participants. These participants are not centrally controlled: they join and leave freely, and can also drop out unexpectedly during training. This implies that autoscaling is triggered by each participating node rather than by a central “GPU scheduler”. As a result:
  • (a)
    Autoscaling occurs more frequently, since node arrivals and departures are no longer batch operations but individual events. This places stricter requirements on the latency of scale-in and scale-out.
  • (b)
    Autoscaling and self-healing can affect model quality. The addition or removal of nodes corresponds to the injection or withdrawal of training data. Thus, the impact of scale-in and scale-out on model quality should be investigated.
This paper proposes Chaos, a multi-party distributed training system with self-healing and autoscaling, enabling robust and elastic training under churn. With this system, institutions can freely join or leave during training and alter their connections. Node or link join, leave, or failure initiates scaling events, then neighbors run peer negotiation protocols to perform scale-in, scale-out, connect-link, and disconnect-link. These protocols remove faulty nodes or links, fix communication rules, add or remove participants, and resume training in a fully self-governed manner. To speed up autoscaling, we propose multi-neighbor state replication with model sharding. With this mechanism, a new node can fetch the latest state in parallel from multiple neighbors while balancing their transmission loads to minimize scale-out delay. We set up a cluster with 6–12 parties, connect them with a constrained, heterogeneous WAN, and simulate node joins and exits. Compared to Pollux [19], Elan [30], and Autoscaling [31], Chaos handles self-healing and autoscaling more efficiently, with scale-out delays between 0.5-5 seconds and others within 20ms. Our main contributions are summarized as follows:
@@
  • We propose a multi-party distributed training system with self-healing and autoscaling that can more efficiently handle node and link joins, exits, and failures.
  • We formulate the multi-neighbor state replication problem with model sharding and assignment, derive a greedy solution, and prove its optimality. This yields a network-aware strategy that minimizes replication time.
@@
  • We propose a cluster monitor that tracks resource and topology changes, along with peer negotiation protocols that enable self-governed autoscaling.
@@
  • We implement Chaos and demonstrate that it achieves substantially lower scale-out delay than Pollux, Elan, and Autoscaling, with high resource efficiency and scalability.
@@ II Related Work and Motivations
In this section, we first review related work on self-healing and autoscaling in distributed machine learning (DML), then explain how our work differs, outline our focus, and present the motivation behind the design of Chaos.
@@ II-A Reviews on Self-healing and Autoscaling
Self-healing. In DML, worker nodes should synchronize model weights frequently, but in chaotic WANs, frequent node and link failures block synchronization and freeze training. Checkpointing helps recover from such failures by periodically saving the training state, so the training can restart from the last checkpoint [13]. DataStates-LLM [14] and CheckFreq [15] reduce training stalls by overlapping computation and checkpoint I/O. JIT Checkpointing [16] creates a remote checkpoint only after failures. Check-N-Run [17] combines differential checkpoints with quantization and decoupled background persistence to cut checkpoint size and bandwidth without hurting accuracy. Gemini [18] uses remote CPU memories and high-speed networks to achieve fast checkpointing.
@@
Autoscaling. Checkpointing is also a common approach to autoscaling, known as “stop-resume[28]: it saves a checkpoint, pauses training, resizes the cluster, then restarts from the latest checkpoint. To reduce the checkpoint-restart overhead, Optimus [21] reconfigures GPU resources at a fixed ten-minute interval. Pollux [19] executes reconfiguration and saves checkpoints when cluster size and topology have to change. DeepBoot [22] uses auto-fast elastic to improve Pollux’s cold restarts into fast checkpoint I/O and light communication reconfiguration, thus reducing rescheduling delay. Rubick [20] performs cost-aware reconfiguration: executes reconfiguration when the restart overhead stays below an empirical threshold. Universal Checkpointing [23] reduces reconfiguration time with parallel I/O, high-bandwidth broadcasting, and lazy execution. Nevertheless, they still take more than one minute.
@@
Stop-free autoscaling offers a smoother alternative. When a new node joins, it fetches the training state from existing nodes’ memory, so there is no need to restart or load/write checkpoints from/to a slow disk. In parameter server setups, PaddleEDL [24] and ElasticDL [25] use asynchronous training to realize stop-free scaling. As workers run independently, their join and exit don’t disrupt others, and new workers fetch the latest model weights from the parameter server. However, they fall back to stop-resume in synchronous mode. DL2 [26] and EPS [27] enable stop-free scaling in synchronous mode, and also require new nodes to fetch the training state from parameter servers. EDL [28] and CoDDL [29], on the other hand, fetch from a single worker node. Autoscaling [31] goes further by slicing the training state across multiple nodes and enabling parallel fetching for faster replication. Elan [30], in contrast, explores a different path by having the new node fetch from the neighbor with the highest available bandwidth.
II-B Differences and Focus of This Paper
The prior works focus on optimizing reconfiguration policies (e.g., adjusting execution plans and GPU allocations) via a centralized GPU scheduler, but deploy new configurations via simple checkpoint-restart or naive stop-free approaches, leading to minute-scale delays during scale-in and scale-out. Instead, Chaos targets minimizing these scale-in and scale-out delays, and in our scenario:
  • A centralized node/GPU scheduler doesn’t exist as scaling events are triggered independently by each self-governed institution, which can join or leave freely. As a result, nodes join and leave one at a time, rather than in batches.
  • During scaling, training data joins or leaves with the institution node, unlike in HPC, where data can be migrated to other nodes. This may affect model quality.
Therefore, cluster scaling will occur more often, demanding lower scaling latency. These differences distinguish Chaos from prior work.
Gossip FL (peer-to-peer, decentralized) also supports self-healing and autoscaling. However, recent studies have shown that gossip protocols often fail to converge under high churn [32], where nodes frequently join and leave, while centralized FL preserves both convergence and accuracy [33]. Therefore, we focus on the centralized setup. Please note that while Chaos uses a centralized architecture for the data plane (e.g., the transmission of model updates), our scaling-delay optimization in the control plane follows a decentralized design, where neighbor institutions can exchange control messages directly. This decoupling of data and control planes makes this work complementary to existing optimizations. For example, we can correct model updates on the server side to restore training stability under high churn [34]; adjust worker workload to mitigate stragglers [35]; schedule data-plane traffic to accelerate model synchronization [36]; and select training data or nodes to reduce non-IID impact [37]. These techniques can be directly applied, but are not the focus of this paper.
@@ II-B Differences and Focus of This Paper
Figure 2: An example of (a) the overlay network topology, (b) single-node replication, (c) multi-node replication, and (d) multi-neighbor replication. Figure (e) compares their state replication delays. The number on each edge indicates the delay incurred to transfer one unit of data through that link. Suppose there are 10 data units to transmit.
II-C Motivations
Question A (QA): From which node(s) should a new node fetch to speed up state replication?
Scale-out delay is the most significant among scaling actions, and in the worst case, it can reach tens of seconds. This delay depends on which node(s) the new node fetches the latest training state from.
In the overlay network topology shown in Fig. 2a, each node represents an institution. Prior work [28, 29, 30] fetches the training state from a single worker node (see Fig. 2b), which may or may not be a neighbor. Following the idea in [30], when a new node 11 joins, it fetches the training state from its highest-bandwidth neighbor 3, but sending the entire state over a single link creates a bottleneck: state replication still takes 80 time units.
Multi-source solutions [24, 25, 26, 27, 31] fetch different shards of the training state in parallel from multiple nodes (Fig. 2c), which can be parameter servers or worker nodes. Even with multi-path transfers, the new node may be far from the sources, requiring multiple hops and sometimes crossing bottleneck links. In Fig. 2c, this can cause delays by up to 88 time units. This approach is used by default in parameter server systems but often suffers from unpredictable efficiency.
This leads to an insight: can a new node fetch different shards of the training state from multiple neighbors? Yes! In synchronous training, parameter servers and workers maintain identical model weights, so a new node can obtain the latest model by fetching at the right time. To handle heterogeneous WAN bandwidth, the model should be appropriately sharded to balance the load across different transfer paths. Fig. 2d shows an example: the new node 11 fetches most model shards from high-bandwidth neighbor 3, while also fetching fewer shards in parallel from lower-bandwidth nodes 4 and 6. In this way, the state replication delay is reduced to 62 time units. This calls for a smart shard scheduler that minimizes the completion time to fetch from neighbors.
Question B (QB): How do nodes join or leave a self-governed cluster?
Prior work [21, 19, 22, 20, 24, 25, 26, 27, 28, 29, 30, 31, 23] has generally relied on a GPU scheduler to make scaling decisions. This central scheduler holds full control over all nodes and GPUs, so it can add or remove them in batches at set times. In Fig. 3a, the admin acts as this scaling scheduler: it sends join commands and authorizes new nodes to participate in the next training round. This paradigm, however, does not apply to self-governed clusters.
Figure 3: Example of node joins with and w/o a central admin.
In a self-governed cluster, institutions join and leave on their own schedule, and no single party controls the others. Thus, a central “scaling scheduler” would lack the authority to command or batch participating parties, and joins and leaves must be handled locally, one at a time, through peer negotiation. As shown in Fig. 3b, each node runs a lightweight agent to discover neighbor and link changes, negotiate admission, trigger scaling actions (e.g., connect-link and scale-out), and synchronize training state to the new node. Similarly, departures or failures are discovered by neighbors, which then negotiate and execute actions (e.g., disconnect-link, scale-in) to restore training. This requires a cluster monitor to detect node and link changes, and peer negotiation protocols to coordinate scale-in, scale-out, connect-link, and disconnect-link actions.
@@ II-C Motivations
To address QA, Section III formulates the model sharding and assignment problem, introduces a polynomial-time algorithm, and analyzes its optimality and complexity. For QB, Section IV presents a cluster monitor and peer negotiation protocols for self-governed scaling. Section V outlines the system design, and Section VI provides experimental results.